Grade 11General Physics

A particle is throw vertically upward. Its velocity at half of the height is 10m/s then maximum height attained by it

Sudipti
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 

let intitial velocity is U and maximum hight is h

so at maximum hight velocity will be zero

0 =U2 -2gh

U2 =2gh ..........1

now given velocity is = 10m/sec at hight h/2

so 102 = U2 -2gh/2

100 =2gh -gh

h = 100/10

h =10 m/sec

 

Regards

Arun (askIITians forum expert)